Integrate x2+x+5/(x2+4)(x+1) dx
1
Expert's answer
2020-10-06T18:34:21-0400

Solution:

"\int \frac{x^2+x+5}{\left(x^2+4\right)\left(x+1\right)}dx"

"=\int \frac{(x^2+4)+(x+1)}{\left(x^2+4\right)\left(x+1\right)}dx"

"=\int (\frac{1}{x^2+4}+\frac{1}{x+1})dx"

"=\int \frac{1}{x^2+4}dx+\int \frac{1}{x+1}dx"

"=\int \frac{1}{x^2+2^2}dx+\int \frac{1}{x+1}dx"

On integrating,

"=\frac{1}{2}\tan^{-1} \left(\frac{x}{2}\right)+\ln \left|x+1\right|+C" [Using "\int \frac{1}{x^2+a^2}dx=\frac{1}{a}\tan^{-1} \left(\frac{x}{a}\right) ;\int \frac{1}{x}dx=\ln|x|" ]

Answer:

"\frac{1}{2}\tan^{-1} \left(\frac{x}{2}\right)+\ln \left|x+1\right|+C"